# Which Similarity-Sensitive Entropy (S-entropy)?

### v2

@@ 4.2 Trends with LCR and VS
When $k$ is very small, all $z_{ij}$ approach 1, collapsing the system to a single individual for both LCR and VS (Fig. 6a and c, left), yielding a perfect correlation ($R^{2}=1.00$). When $k$ is very large, all off-diagonal $z_{ij}$ approach 0, making every element essentially unique; consequently, both LCR and VS converge to the number of elements in the dataset, and the correlation again is perfect ($R^{2}=1.00$) (Fig. 6a and c, right). Only at intermediate $k$ values do LCR and VS provide independent information (Fig. 6a and c, middle).
@@ 4.3 Bounds on LCR and VS
Note that in the context of this conjecture, since we are talking about the LCR diversity on a set with the uniform distribution, there is no distinction between $Z$ and $Z_{n}$ (like that made in section 2.2.2), as these are equal. We prove this for $q=2,3,$ and $\infty$ in Appendix A.2. We consider these proofs strong evidence that Conjecture 1 holds for all $q\geq 2$, since it would be surprising for VS to bound LCR at $q=2,3,$ and $\infty$ but not $q=2.5,4$, or $27.3$. The following lemma supports this conjecture holding for all $q$:
@@
A computational search of $\approx 4,000$ examples involving similarity matrices of up to size $50\times 50$ failed to find counterexamples for any of several values of $q$, including $q=0$ and $q=1$. This search used a method that successfully found a counterexample to a different inequality previously [Couch:2023pav], adding further support for Conjecture 1 being true. We present an approach to a proof for all $q$ in Appendix A.3 and present an additional proof on bounds in Appendix A.4.
@@ 5.1 Richer measures of the information that ML datasets contain
One way to get at this missing information is to use clustering algorithms, to quantify dataset substructure as the number of clusters and possibly clusters’ relative sizes (which can be measured by traditional entropy). Like LCR and VS, which require a choice of similarity measure and $q$, clustering algorithms require choices. For example, HDBSCAN requires choosing min_cluster_size, min_samples, etc. Conceptually, the main difference between clustering and S-entropy is that S-entropy accounts for similarities both within and between clusters, whereas most clustering algorithms (and classifiers) lack this ability: effectively, the within-cluster (or intra-class) similarity is always 1 and the between-cluster similarity is always 0.
@@
The exception is fuzzy clustering, in which each element can belong to more than one cluster, with each element being parameterized by its strength of membership for one or more clusters. One can view LCR as carrying fuzzy-clustering to its logical limit, with each unique element being its own cluster and the similarities giving other elements’ strength of membership to that cluster and thereby contributing to the size of that cluster. This is a fuzzy-clustering interpretation of the ordinariness, $Z\mathbf{p}_{i}$: a big cluster that many unique elements belong to fairly strongly is “ordinary” (Fig. 2c). This perspective provides another way of seeing that LCR is the entropy of a system after accounting/adjusting/correcting for the similarity among its elements.
@@
Traditional entropy accounts for frequency; LCR accounts for both frequency and similarity. In this way, LCR is a richer, more accurate measure of the information in a system. Note that, regardless of whether they are represented as effective numbers (as we [nguyen2023greylock] and others [hill1973diversity, jost2006entropy, leinsterEntropyDiversityAxiomatic2020] have advocated) or as bits of information, LCR is always bounded above by the corresponding traditional entropy [leinsterEntropyDiversityAxiomatic2020]; for example, at $q=1$ LCR never exceeds Shannon entropy. Similarity saps individual elements of some of their uniqueness, but in doing so, it reveals differences between datasets that traditional entropy cannot (Fig. 1).
@@
How do LCR and VS compare with the number of clusters identified by HDBSCAN? Predicting the relationship is difficult: large clusters whose peripheral elements are relatively distinct tend to increase the LCR/VS values, whereas strong similarities among elements—both within a cluster and across neighboring clusters—tend to reduce them (see Figs. 1, 5c). Indeed, many datasets examined exhibited a pronounced difference in magnitude between tabular and imaging datasets when comparing the HDBSCAN cluster count with LCR and VS. For tabular data sets, both LCR and VS almost always surpassed the HDBSCAN cluster count, often by one to two orders of magnitude, reflecting that many rows remain unclustered by hard clustering.
@@
In contrast, for imaging datasets, LCR and VS were typically one to two orders of magnitude lower than the number of clusters, indicating substantial similarity among clusters that hard clustering does not capture (Fig. 2c). We interpret these systematic differences as evidence of the advantages of similarity‑sensitive entropy relative to hard clustering, for describing a system’s structure. We do not regard this as a failure of HDBSCAN; we note as a limitation of our study that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity‑sensitive entropy and clustering. However, we note that by the “eye test,” UMAP plots validate HDBSCAN’s clustering, suggesting hyperparameter search would be unlikely to meaningfully affect our findings.
@@ 5.2 The effect of scaling the similarity measure/similarity matrix
We demonstrated that the relationship between LCR and VS depends strongly on how the similarity‑ matrix values are scaled, which we illustrated by introducing the half‑distance parameter $k$. Recall that similarity between two elements ranges from 0 to 1; scaling modifies this relationship by raising the values to a power, thereby shifting the curve upward or downward in the middle, while maintaining the anchoring at each end (at 0 and 1).
Why should one need to scale the similarity between elements in the first place? Should scaling be permitted at all? When similarity is derived from external knowledge or some principle or outside requirement, we argue that scaling should not be applied: the similarity function is defined a priori, its output for any given pair is fixed, and the similarity matrix simply collects these values for all pairs. An example is binding similarity in immune repertoires, defined as the relative dissociation constants of two antibodies (or TCRs) for a given antigen (or set of antigens) [aroraRepertoirescaleMeasuresAntigen2022]. This is the unique definition that preserves biophysical additivity; arbitrary scaling would break this relationship and thereby produce a similarity matrix that is objectively incorrect.
In contrast, when similarity is heuristic and not constrained by such principles, as with the Euclidean‑distance-based similarity measures used here, there is no a priori reason to privilege the default similarity ($k=1$) over other values of $k$. For instance, RMSD provides a perfectly valid similarity definition and corresponds to Euclidean distance with $k=1/\sqrt{d}$, where $d$ is the dimension of the underlying data (number of columns for tabular data, number of pixels $\times$ number of color channels for image data, 2 for the umap embedding, etc). Therefore, in heuristic cases, scaling can serve as a reasonable dial that investigators can adjust to obtain values that make the most sense given the context and/or are maximally informative about the system.
Such choices arise often in measurement. For example, we note that the parameter $q$ offers a loosely analogous scaling to that used in traditional (and similarity-sensitive) entropy, allowing investigators to adjust the emphasis on frequency. Lest choices feel too arbitrary, remember that specific integer values of $q$ correspond to well‑known statistics; for example, $q=2$ corresponds to Simpson’s index, which reflects the statistics of collision frequency. It is possible that alternative values of $k$ could similarly be linked to other quantities of interest. Investigating this possibility is left for future work.
@@
So: for heuristic similarity functions, which $k$ should one choose? Qualitatively, the answer is a value roughly midway between the extremes (Fig. 6). However, our quantitative findings indicate that the optimal choice is not obvious. We found that the correlation between LCR and VS is minimized at $k=0.2$ for tabular data sets and $3.4$ for imaging data sets—not at any obvious definition of the midpoint of the range (defined for example as the geometric mean of the $k$ values where the correlation first falls to 0.99 on each side). Moreover, the shapes of these curves, with their multiple extrema, do not appear to follow a simple parametric form, making it difficult to predict precisely where the two metrics are maximally mutually informative. We also observed that even at the $k$ that minimizes the LCR-VS correlation, the $R^{2}$ between LCR and VS remained relatively high in absolute terms, at $0.68$. Perhaps this was to be expected, since both metrics quantify the effective number of elements using the same underlying similarity measure, albeit in different ways (see Section 6). The dependence on $k$, however, was quite strong.
Our results suggest first testing whether the entropy lies near 1 or $n$; if so, to explore, if possible, different $k$ values to delineate a non‑trivial range; and finally to select a $k$ approximately in the middle of this range (on a logarithmic scale). Both the choice of similarity measure and the value of $k$ have effects on the resulting entropy that cannot be ignored. Thus to some extent, the information in a system is in the eye of the beholder.
@@ 6 Conclusion
From a practical perspective, LCR has three advantages. First, it is less computationally expensive whenever elements appear multiple times (i.e., when $\mathbf{p}$ is not flat/uniform), because the similarity matrix for LCR has order equal to the number of unique elements, whereas VS requires a matrix whose order equals the total number of elements; both calculations scale as the square of the matrix order. Empirically, we observed a speed advantage of greylock over vendi-score even when all elements were unique, a situation that frequently occurs in ML datasets (Fig. 8). This performance gap may narrow or widen as codebases evolve.
@@
Figure 8: Compute times for all 53 datasets for LCR (using the greylock Python package [nguyen2023greylock]) vs. VS (using vendi-score). Imaging datasets were all 10,000 images. Smaller datasets were faster. Note the different timescales.
@@
Second, LCR does not require the similarity matrix to be PSD, whereas VS does. Similarity matrices need not be PSD; a perfectly reasonable set of similarity measurements can produce a matrix that fails the PSD condition, which would complicate the use of VS. Even symmetric similarity matrices—the most common and most natural type, satisfying $z_{ij}=z_{ji}$, so that similarity is reciprocal—are not guaranteed to be PSD, even when their entries obey the triangle inequality. The prevalence of such non‑PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD (a consequence of how we chose our similarity measures; this follows from Theorem 2.11 in [leinster2013magnitude]). Thus, PSD‑ness is irrelevant for LCR but must be checked to apply VS.
@@
And third, LCR in its various forms [chao2010phylogenetic, chaoAttributediversityApproachFunctional2019] (Table 1) is part of an extensively characterized framework [leinsterEntropyDiversityAxiomatic2020] that provides similarity‑sensitive analogs of mutual information, relative entropy, and cross‑entropy and offers methods for partitioning the overall similarity‑sensitive entropy ($\gamma$) into within‑class ($\alpha$) and between‑class ($\beta$) components. In entropic terms, $\alpha$ is the (effective‑number form of the) joint entropy of the elements together with their class label; $\beta$ is the mutual information between the elements and the class label; and $\gamma$ is the marginal entropy of the elements irrespective of class. In this work we computed LCR in its $\gamma$ form; when only a single class is present, $\gamma$ coincides with $\alpha$. Measuring the full $\alpha$, $\beta$, and $\gamma$ decomposition across ML classes is a promising direction for future research. (The greylock implementation computes $\alpha$, $\beta$, and $\gamma$.)
@@
The necessary mathematics for extending this decomposition to VS exists, because VS is based on von Neumann (quantum) entropy, for which quantum mutual information, quantum relative entropy, and related quantum quantities are well understood [nielsen2010quantum]. However, computing these quantities requires a matrix that simultaneously encodes each element and its class membership. If there are $n$ total elements and $k$ classes, the VS decomposition demands either an $nk\times nk$ matrix or an $n\times k\times n\times k$ 4-dimensional tensor (which can be reduced to $n\times n$ and $k\times k$ matrices by taking a partial trace). In contrast, LCR requires only the similarity matrix $Z$. Consequently, VS needs an explicit notion of similarity both between classes and between elements, whereas in LCR any class‑level similarity arises purely from element‑level similarities within the classes. In principle, the VS formulation would permit the confusing situation in which two classes contain exactly the same individuals but nonetheless have zero similarity—a scenario that cannot occur under LCR, in which the similarity between classes is purely a function of similarities among the elements they contain.
@@
In sum, LCR and VS provide complementary descriptions of similarity‑sensitive entropy, a unifying concept that is likely to find applications in many domains in which traditional entropy is currently a mainstay. We anticipate that the deeper understanding of their relationship and properties through work such as that presented here will aid such investigations.

### v3

@@ 4.2 Trends with LCR and VS
The half-distance $k$ has a major impact on the correlation between LCR and VS. When $k$ is very small, all $z_{ij}$ approach 1, collapsing the system to a single individual for both LCR and VS (Fig. 6a and c, left), yielding a perfect correlation ($R^{2}=1.00$). When $k$ is very large, all off-diagonal $z_{ij}$ approach 0, making every element essentially unique; consequently, both LCR and VS converge to the number of elements in the dataset, and the correlation again is perfect ($R^{2}=1.00$) (Fig. 6a and c, right). Only at intermediate $k$ values do LCR and VS provide independent information (Figs. 6a and 6c, middle).
@@ 4.3 Bounds on LCR and VS
Note that in the context of this conjecture, since we are considering the LCR diversity on a set with the uniform distribution, there is no distinction between $Z$ and $Z_{n}$ (like that made in section 2.2.2), as these are equal. We prove this conjecture for $q=2,3,$ and $\infty$ in Appendix A.2. We consider the existence of proofs for these several $q$ to be strong evidence that Conjecture 1 holds for all $q\geq 2$, since it would be surprising for VS to bound LCR at $q=2,3,$ and $\infty$ but not, for example, at $q=2.5,4$, or $27.3$. The following lemma supports this conjecture holding for all $q$:
@@
A computational search of $\approx 4,000$ examples involving similarity matrices of up to size $50\times 50$ failed to find counterexamples for any of several values of $q$, including $q=0$ and $q=1$. The method behind this search has been successfully used to found counterexamples to other inequalities [8], adding further support for Conjecture 1 being true. We present an approach to a proof for all $q$ in Appendix A.3 and present an additional proof on bounds in Appendix A.4.
@@ 5.1 Richer measures of the information that ML datasets contain
One way to get at this missing information is to use clustering algorithms to quantify dataset substructure as the number of clusters and possibly clusters’ relative sizes (which can be measured by traditional entropy). Like LCR and VS, which require a choice of similarity measure and $q$, clustering algorithms require choices about parameters; for example, HDBSCAN requires choosing min_cluster_size, min_samples, etc.
@@
Conceptually, the main difference between clustering and S-entropy is that S-entropy accounts for similarities both within and between clusters, whereas most clustering algorithms (and classifiers) lack this ability: effectively, the within-cluster (or intra-class) similarity is always 1 and the between-cluster similarity is always 0. The exception is fuzzy clustering, in which each element can belong to more than one cluster, with each element being parameterized by its strength of membership for one or more clusters. One can view LCR as carrying fuzzy-clustering to its logical limit, with each unique element being its own cluster and the similarities giving other elements’ strength of membership in that cluster and thereby contributing to the size of the cluster. This is a fuzzy-clustering interpretation of the ordinariness, $Z\mathbf{p}_{i}$: a big cluster to which many unique elements belong fairly strongly would be considered “ordinary” (Fig. 2c). This perspective provides another way of seeing that LCR is the entropy of a system after accounting/adjusting/correcting for the similarity among its elements.
@@
Traditional entropy accounts for frequency; LCR accounts for both frequency and similarity. In this way, LCR is a richer, more complete measure of the information in a system. Note that, regardless of whether they are represented as effective numbers (as we [23] and others [12, 13, 17] have advocated) or as bits of information, LCR is always bounded above by the corresponding traditional entropy [17]; for example, at $q=1$ LCR never exceeds Shannon entropy. Thus, similarity can be seen as sapping individual elements of some of their uniqueness. However, in doing so, it reveals differences between datasets that traditional entropy cannot (Fig. 1).
@@
How do LCR and VS compare with the number of clusters identified by HDBSCAN? Predicting the relationship is difficult: large clusters whose peripheral elements are relatively distinct tend to increase the LCR/VS values, whereas strong similarities among elements—both within a cluster and across neighboring clusters—tend to reduce them (see Figs. 1, 5c). Indeed, many of the datasets we examined exhibited a pronounced difference in magnitude between tabular and imaging datasets when comparing the HDBSCAN cluster count with LCR and VS. For tabular data sets, both LCR and VS almost always surpassed the HDBSCAN cluster count, often by 1-2 orders of magnitude, reflecting that many rows remain unclustered by hard clustering (these do not contribute the the HDBSCAN total).
@@
In contrast, for imaging datasets, LCR and VS were typically 1-2 orders of magnitude lower than the number of clusters, indicating substantial similarity among elements across clusters, which hard clustering does not capture (Fig. 2c). We interpret these systematic differences as evidence of the advantages of S-entropy relative to hard clustering, for describing a system’s structure. We do not regard this as a failure of HDBSCAN; a limitation of our study that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity‑sensitive entropy and clustering. However, we note that by the “eye test,” UMAP plots validate HDBSCAN’s clustering, suggesting hyperparameter search is unlikely to meaningfully affect the findings or conclusions presented.
@@ 5.2 The effect of scaling the similarity measure/similarity matrix
We demonstrated that the relationship between LCR and VS depends strongly on how the similarity‑ matrix values are scaled, which we illustrated by introducing the half‑distance parameter $k$. Recall that similarity between two elements ranges from 0 to 1; scaling modifies this relationship by raising the values to a power, thereby shifting the curve in the middle of this range, while maintaining the anchoring at each end (at 0 and 1).
Why should one need to scale the similarity between elements in the first place? Should scaling be permitted at all? When similarity is derived from external knowledge or some principle or outside requirement, we argue that scaling should not be applied: the external knowledge defines and thereby constrains the similarity function a priori, its output for any given pair is fixed, and the similarity matrix simply collects these values for all pairs. An example is binding similarity in immune repertoires, defined as the relative dissociation constants of two antibodies (or TCRs) for a given antigen (or set of antigens) [2]: this is the unique definition that preserves biophysical additivity; arbitrary scaling would break this relationship and thereby produce a similarity matrix that is objectively incorrect.
In contrast, when similarity is heuristic and not constrained by such knowledge, as with the Euclidean‑distance-based similarity measures used here, there is no a priori reason to privilege the default similarity ($k=1$) over other values of $k$. For instance, RMSD provides a perfectly valid similarity definition and corresponds to Euclidean distance with $k=1/\sqrt{d}$, where $d$ is the dimension of the underlying data (number of columns for tabular data, number of pixels $\times$ number of color channels for image data, 2 for the umap embedding, etc). Therefore, in heuristic cases, scaling can serve as a reasonable dial that investigators can adjust to obtain values that make the most sense given the context/are maximally informative about the system. Such choices arise often in measurement. For example, the parameter $q$ is an analogous dial in traditional entropy (as well as S-entropy), allowing investigators to adjust the emphasis on frequency. As with $k$, prior information can constrain choice of $q$; for example, when the measurement has to do with collision frequency, $q=2$ should be used, consistent with $q=2$ corresponding to Simpson’s index. Whether specific values of $k$ might similarly be linked to known quantities of interest is left for future work.
@@
For heuristic similarity functions, which $k$ should one choose? Qualitatively, the answer is a value roughly midway between the extremes (Fig. 6). However, our quantitative findings indicate that the optimal choice is not obvious. We found that the correlation between LCR and VS is minimized at $k=0.2$ for tabular data sets and $3.4$ for imaging data sets—not at any obvious definition of the midpoint of the range (defined for example as the geometric mean of the $k$ values where the correlation first falls to 0.99 on each side). Moreover, the shapes of these curves, with their multiple extrema, do not appear to follow a simple parametric form, making it difficult to predict precisely where the two metrics are maximally mutually informative. We also observed that even at the $k$ that minimizes the LCR-VS correlation, the $R^{2}$ between LCR and VS remained relatively high in absolute terms, at $0.68$. Perhaps this was to be expected, since both metrics quantify the effective number of elements using the same underlying similarity measure, albeit in different ways (see Section 6). The dependence on $k$, however, was quite strong.
Our results suggest first testing whether the entropy lies near 1 or $n$; if so, to explore, if possible, different $k$ values to delineate a non‑trivial range; and finally to select a $k$ approximately in the middle of this range (on a logarithmic scale). Both the choice of similarity measure and the value of $k$ have effects on the resulting entropy that cannot be ignored. Thus to some extent, from a quantitative perspective, the information in a system is in the eye of the beholder.
@@ 6 Conclusion
From a practical perspective, LCR has three advantages. First, it is less computationally expensive whenever elements appear multiple times (i.e., when $\mathbf{p}$ is not flat/uniform), because the similarity matrix for LCR has order equal to the number of unique elements, whereas VS requires a matrix whose order equals the total number of elements; both calculations scale as the square of the matrix order. Empirically, we observed a speed advantage of sentropy over vendi-score even when all elements were unique, a situation that frequently occurs in ML datasets (Fig. 8).
@@
Figure 8: Compute times for all 53 datasets for LCR (using the sentropy Python package [23]) vs. VS (using vendi-score). Imaging datasets (blue) were all 10,000 images. Smaller datasets were faster. Note the different timescales.
@@
Second, LCR does not require the similarity matrix to be PSD, whereas VS does. Similarity matrices need not be PSD; a perfectly reasonable set of similarity measurements can produce a matrix that fails the PSD condition, which would complicate the use of VS. Even symmetric similarity matrices—the most common and most natural type, satisfying $z_{ij}=z_{ji}$, so that similarity is reciprocal—are not guaranteed to be PSD, even when their entries obey the triangle inequality. The prevalence of such non‑PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD (a consequence of how we chose our similarity measures; this follows from Theorem 2.11 in [16]). Thus, PSD‑ness is irrelevant for LCR but must be checked to apply VS.
@@
And third, LCR in its various forms [3, 4] (Table 1) is part of an extensively characterized framework [17] that provides similarity‑sensitive analogs of mutual information, relative entropy, and cross‑entropy and offers methods for partitioning the overall similarity‑sensitive entropy ($\gamma$) into within‑class ($\alpha$) and between‑class ($\beta$) components. In entropic terms, $\alpha$ is the (effective‑number form of the) joint entropy of the elements together with their class label; $\beta$ is the mutual information between the elements and the class label; and $\gamma$ is the marginal entropy of the elements irrespective of class. In this work we computed LCR in its $\gamma$ form; when only a single class is present, $\gamma$ coincides with $\alpha$. Measuring the full $\alpha$, $\beta$, and $\gamma$ decomposition across ML classes is a promising direction for future research. (The sentropy package computes $\alpha$, $\beta$, and $\gamma$.)
@@
The necessary mathematics for extending this decomposition to VS exists, because VS is based on von Neumann (quantum) entropy, for which quantum mutual information, quantum relative entropy, and related quantum quantities are well understood [24]. However, computing these quantities requires a matrix that simultaneously encodes each element and its class membership. If there are $n$ total elements and $k$ classes, the VS decomposition demands either an $nk\times nk$ matrix or an $n\times k\times n\times k$ 4-dimensional tensor (which can be reduced to $n\times n$ and $k\times k$ matrices by taking a partial trace). In contrast, LCR requires only the similarity matrix $Z$. Consequently, VS needs an explicit notion of similarity both between classes and between elements, whereas in LCR any class‑level similarity arises purely from element‑level similarities within the classes. In principle, the VS formulation would permit the confusing situation in which two classes contain exactly the same individuals but nonetheless have zero similarity—a scenario that cannot occur under LCR, in which the similarity between classes is purely a function of similarities among the elements they contain.
@@
In sum, LCR and VS provide complementary descriptions of similarity‑sensitive entropy, a unifying concept that is likely to find applications in many domains in which traditional entropy is currently a mainstay. We anticipate that the deeper understanding of their relationship and properties will aid such investigations.